# Deep Learning Approach for Changepoint Detection: Penalty Parameter Optimization

### v1

@@ 1 Introduction
Previous methods, such as those employing Bayesian Information Criterion (BIC) (Schwarz, 1978), linear models (Hocking and Srivastava, 2023; Rigaill et al., 2013), and Maximum Margin Interval Trees (MMIT) (Drouin et al., 2017) have made good attempts to ascertain the optimal $\lambda$ value. However, these methods rely on using linear models or decision trees as learning models may constrain the ability to capture complex patterns.
@@ 2.1 Problem setting
Changepoint detection algorithms take into the sequence $\mathbf{d}$ along with the penalty parameter $\lambda$ as input and then generate the set of detected changepoints. From each sequence $\mathbf{d}$, $M$ distinct labels $\{(s_{i},e_{i},c_{i})\}_{i=1}^{M}$ emerge within the sequence, where $c_{i}$ denotes the expected number of changepoints between points at locations $s_{i}$ and $e_{i}$. There are two types of label errors: in each label, if the number of detected changepoints greater than $c_{i}$, it’s a false positive; if smaller, it’s a false negative. Finding the optimal set of detected changepoints require considering a range of values for the penalty parameter $\lambda$ (rather than just one specific value) that minimizes the total number of label errors. Consequently, the value of penalty parameter $\lambda$ to be predicted should also fall within an optimal interval, see Figure 2.
@@ 4 Experiments
n layer n neurons detailed (N=3730) systematic (N=3418) epigenomic (N=4913) total
mlp.1 mlp.2 mlp.4 mlp.117 mlp.1 mlp.2 mlp.4 mlp.117 mlp.1 mlp.2 mlp.4 mlp.117
1 $\leq 64$ 0 1 0 1 1 1 2 1 0 2 0 2 11
$>64$ 0 0 0 0 2 0 1 1 0 0 0 0 4
2 $\leq 64$ 1 2 4 1 0 0 0 2 2 1 5 4 22
$>64$ 0 0 0 2 0 0 1 0 0 0 0 0 3
3 $\leq 64$ 5 2 0 0 1 1 1 0 4 3 0 0 17
$>64$ 0 0 0 0 0 2 0 0 0 0 0 0 2
4 $\leq 64$ 0 1 1 2 2 1 0 1 0 0 1 0 9
$>64$ 0 0 1 0 0 1 1 1 0 0 0 0 4
For each model MLP from each dataset, 6 folds are tested, resulting in the selection of 6 MLP configurations. Observing the trend in the table, it’s evident that the preference leans towards MLPs with 2 or 3 hidden layers and the number of neurons is less than 64.

### v2

@@ 1 Introduction
Previous methods, such as those employing Bayesian Information Criterion (BIC) (Schwarz, 1978), linear models (Hocking and Srivastava, 2023; Rigaill et al., 2013), ALPIN (Truong et al., 2017) and Maximum Margin Interval Trees (MMIT) (Drouin et al., 2017) have made good attempts to ascertain the optimal $\lambda$ value. However, these methods rely on using linear models or decision trees as learning models may constrain the ability to capture complex patterns.
@@ 2.1 Problem setting
Changepoint detection algorithms take into the sequence $\mathbf{d}$ along with the penalty parameter $\lambda$ as input and then generate the set of detected changepoints. From each sequence $\mathbf{d}$, $M$ distinct labels $\{(s_{i},e_{i},c_{i})\}_{i=1}^{M}$ emerge within the sequence, where $c_{i}$ denotes the expected number of changepoints between points at locations $s_{i}$ and $e_{i}$. Since ALPIN (Truong et al., 2017) requires a predefined expert sequence partition (or equivalently, each positive label having a size of 1, which is a specific special case), it is not suitable for this problem setting. There are two types of label errors: in each label, if the number of detected changepoints greater than $c_{i}$, it’s a false positive; if smaller, it’s a false negative. Finding the optimal set of detected changepoints require considering a range of values for the penalty parameter $\lambda$ (rather than just one specific value) that minimizes the total number of label errors. Consequently, the value of penalty parameter $\lambda$ to be predicted should also fall within an optimal interval, see Figure 2.
@@ 4 Experiments
layers neurons detailed systematic epigenomic total
1 2 4 117 1 2 4 117 1 2 4 117
1 $\leq 64$ 1 1 1 1 2 1 2 2 11
$>64$ 2 1 1 4
2 $\leq 64$ 1 2 4 1 2 2 1 5 4 22
$>64$ 2 1 3
3 $\leq 64$ 5 2 1 1 1 4 3 17
$>64$ 2 2
4 $\leq 64$ 1 1 2 2 1 1 1 9
$>64$ 1 1 1 1 4
For each model MLP from each dataset, 6 folds are tested, resulting in the selection of 6 MLP configurations. Dataset detailed has 3,730 sequences, systematic has 3,418, and epigenomic has 4,913. Observing the trend in the table, it’s evident that the preference leans towards MLPs with 2 or 3 hidden layers and the number of neurons is less than 64. Specifically, out of the 72 chosen architectures, 15 have 1 hidden layer, 25 have 2 hidden layers, 19 have 3 hidden layers, and 13 have 4 hidden layers.